OpenStudy (anonymous):

Look at the net of the given polyhedron. The sum of the number of edges and vertices of the polyhedron is ________.

OpenStudy (anonymous):

where's the net?

OpenStudy (anonymous):

An edge is one of the straight lines, and the vertices are where the lines meet. You just have to count them.

OpenStudy (btaylor):

By Euler's formula (F + V - E = 2): There are 8 faces, so F = 8 There are 8 triangles, and each has three edges. So E=24 8+V-24=2 Solve for V: V=18. So E+V=24+18=42

OpenStudy (anonymous):

ok thank-you guys!!